# Yong Ji1, Junye Li1 and Rui Yang∗2,3,4

### v2

@@ 3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
Recall that $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$. By specification property, there exists $y_{k,j}\in X$ so that the pieces of orbits $\{x_{k,j},\cdots,f^{n_{(k,j)-1}}(x_{(k,j)})\}$ repeating $N_{(k,j)}$ times, can be $\epsilon_{(k,j)}$-traced by $y_{(k,j)}$ with gap $m_{(k,j)}$. Now set $\mathcal{S}_{(k,j)}=\{y_{k,j}\}$, and $c_{k,j}=n_{(k,j)}N_{(k,j)}+(N_{(k,j)}-1)m_{(k,j)}$.
@@
For any $(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}$, let $y(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in X$ be given by the specification property, such that the orbit $\epsilon_{(k,Q_{k})}$-shadows, with gap $m_{(k,Q_{k})}$, the pieces of orbits $\{y_{i_{l}}^{k},f(y_{i_{l}}^{k}),\ldots,f^{n_{k}-1}(y_{i_{l}}^{k})\}$, $l=1,2,\ldots,N_{(k,Q_{k})}$. Then define
$\mathcal{S}_{(k,Q_{k})}=\left\{y=y(i_{1},\ldots,i_{N_{(k,Q_{k})}}):(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}\right\}.$
Write $c_{k,Q_{k}}=n_{k}\cdot N_{(k,Q_{k})}+(N_{(k,Q_{k})}-1)m_{(k,Q_{k})}$, then it follows from the facts that $\mathcal{C}_{(k,Q_{k})}$ is a $(n_{k},4\epsilon^{\ast})$-separated set, and $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set. In particular, by (3.6) and (3.8), one has
@@
Similar to the argument in (3.7) one has
$D(\mathcal{E}_{c_{k,j}}(y_{k,j}),\mu_{(k,j)})<\frac{2}{k}+\delta_{(k,j)},\ y_{k,j}\in\mathcal{S}_{(k,j)}\ \text{for }k\geq 1\ \text{and }1\leq j<Q_{k},$
@@
Hence, in an unified way we always have
@@
Step $2$. Constructions of $\mathbb{F}_{(k,j)}$ and Moran set $\mathbb{F}$.
Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$. Now suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$, and then we will describe how to construct $\mathcal{I}_{(k,j)}$. Let
@@
#$\#\mathcal{S}_{(k,Q_{k})}\geq\exp\left[(M_{(k,Q_{k})}-M_{(k-1,Q_{k-1})})(1-\zeta)^{2}(H-3\zeta)\right].$
For $x\in\mathcal{I}_{(k,j)^{\ast}}$ and $y\in\mathcal{S}_{(k,j)}$, by specification property let $z=z(x,y)\in X$ be some point such that
@@
We set $\mathcal{I}_{(k,j)}=\{z(x,y):x\in\mathcal{I}_{(k,j)^{\ast}},y\in\mathcal{S}_{(k,j)}\}$. For $x\in\mathcal{I}_{(k,Q_{k})^{\ast}}$ and distinct $y,y^{\prime}\in\mathcal{S}_{(k,Q_{k})}$, since $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set, one has
$d_{M_{k,Q_{k}}}(z(x,y),z(x,y^{\prime}))>2\epsilon^{\ast},$
@@
#$\#\mathcal{I}_{(k,j)}=\#\mathcal{I}_{(k,j)^{\ast}}\cdot\#\mathcal{S}_{(k,j)}=\Pi_{(k^{\prime},j^{\prime})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},j^{\prime})}=\Pi_{(k^{\prime},Q_{k^{\prime}})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}.$
@@
It is easy to see that for each $(k,j)\in\mathcal{A}$ the following is satisfied:
@@
$(2)$ if $z\in\mathcal{I}_{(k,j)}$ descends from $x\in\mathcal{I}_{(k,j)^{\ast}}$, one has $\overline{B}_{M_{(k,j)}}(z,2\epsilon_{(k,j)})\subset\overline{B}_{M_{(k,j)^{\ast}}}(x,2\epsilon_{(k,j)^{\ast}})$. That is $\{\mathbb{F}_{(k,j)}:(k,j)\in\mathcal{A}\}$ is a sequence of decreasing nonempty closed subsets of $X$.
Let
@@
One has the fractal set $\mathbb{F}$ is contained in $\mathbb{F}\subset G_{K}^{\omega,U}$, that is,
@@
Proof.
Firstly, we know $\mathbb{F}\subset\mathbb{F}_{(0,0)}=\overline{B}_{n(0,0)}(a_{0},\frac{\epsilon^{\ast}}{8})\subset\overline{B(a_{0},\tilde{\epsilon})}\subset U$. For $p\in\mathbb{F}$ and any $k\geq 1$, we have $p\in\mathbb{F}_{(k,0)}$. Hence for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there is a $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<3\epsilon_{(k,0)}$, which implies that the orbit of $p$ is $4\epsilon_{(k,0)}$-dense in $X$. By the arbitrariness of $k\geq 1$ and $p\in\mathbb{F}$, we have $\mathbb{F}\subset\{x\in X:\omega_{f}(x)=X\}$.
For $p\in\mathbb{F}$ and sufficiently large $n\in\mathbb{N}$, in order to estimate $D(\mathcal{E}_{n}(p),K)$, we split into four cases to discuss.
@@
Case $(2)$. There exists $(k,j)\in\mathcal{A}$ with $j\neq 0$, such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
@@
Case $(3)$. There exist $(k,1)\in\mathcal{A}$ and $i>1$, such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
@@
Case $(4)$. There exists $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$, such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
@@
To sum up, whatever case we consider, by letting $n\to\infty$ we get $V_{f}(p)\subset K$. Based on above arguments, one can also show sequences $\{\mathcal{E}_{n}(p)\}_{n\in\mathbb{N}}$ and $\{\mu_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ have the same limit points. Notice that $K=\overline{\{\mu_{(k^{\prime},j^{\prime})}:(k^{\prime},j^{\prime})>(k,j)\}}$ holds for any fixed $(k,j)\in\mathcal{A}$. Consequently, we have $V_{f}(p)=K$, and hence $\mathbb{F}\subset G_{K}$. ∎
Step $3$. Construction of measure.
Now we define a probability measure $\rho$ on $\mathbb{F}$ satisfying that the $\rho$-measure of some appropriate dynamical balls decay exponentially fast along sequence $\{M_{k}\}$, where $M_{k}=M_{k,Q_{k}}$. Then we can use the Entropy Distribution Principle developed in Lemma 2.2 to estimate the lower bound of $\epsilon$-packing topological entropy of $\mathbb{F}$.
@@
For simplicity, we write $\mathcal{I}_{k}$, $\mathbb{F}_{k}$ instead of $\mathcal{I}_{(k,Q_{k})}$, $\mathbb{F}_{(k,Q_{k})}$, respectively. Since $M(X)$ is compact, we may assume that $\rho_{k}\to\rho\in M(X)$ as $k\to\infty$. It is clear that $\rho_{k}(\mathbb{F}_{k})=1$ for every $k$, and for any $k>k_{0}$, one has $\rho_{k}(\mathbb{F}_{k_{0}})=1$ since $\mathbb{F}_{k}\subset\mathbb{F}_{k_{0}}$. Thus $\rho(\mathbb{F}_{k_{0}})\geq\limsup_{k\to\infty}\rho_{k}(\mathbb{F}_{k_{0}})=1$, which implies that $\rho(\mathbb{F})=1$. By (3.9), we have
#$\#\mathcal{I}_{k}\geq\exp\left[M_{k}\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
@@
Let $k_{0}\geq 1$, and $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ be any Bowen ball which intersects $\mathbb{F}$. Then for any $k>k_{0}$, one has
@@
Proof.
If $k=k_{0}+1$, it is easy to check that
@@
$\rho_{k}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{\Pi_{k_{0}+1\leq k^{\prime}\leq k}\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}}{\#\mathcal{I}_{k}}=\frac{1}{\#\mathcal{I}_{k_{0}}}.$
For large enough $k_{0}$ and any Bowen ball $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ with $z\in\mathbb{F}$, since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ is open, one has
@@
$\leq\exp\left[-M_{k_{0}}(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
@@
$(1-\zeta)^{2}\cdot(H-3\zeta)\leq h_{top}^{P}(f,\mathbb{F},\frac{\epsilon^{\ast}}{12})\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$
@@
Now dividing $|\log\epsilon^{\ast}|$ in both sides and then taking $\limsup_{\epsilon^{\ast}\to 0}$, we have
@@
$\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$

### v3

@@ 3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
Recall that $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$ is defined above. By specification property again, there exists $y_{k,j}\in X$ so that the pieces of orbits $\{x_{k,j},\cdots,f^{n_{(k,j)-1}}(x_{(k,j)})\}$ repeating $N_{(k,j)}$ times, can be $\epsilon_{(k,j)}$-traced by $y_{(k,j)}$ with the gap $m_{(k,j)}$. Now we set $\mathcal{S}_{(k,j)}=\{y_{k,j}\}$, and $c_{k,j}=n_{(k,j)}N_{(k,j)}+(N_{(k,j)}-1)m_{(k,j)}$.
@@
For any $(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}$, let $y(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in X$ be given by the specification property, such that the orbit $\epsilon_{(k,Q_{k})}$-shadows, with the gap $m_{(k,Q_{k})}$, the pieces of orbits $\{y_{i_{l}}^{k},f(y_{i_{l}}^{k}),\ldots,f^{n_{k}-1}(y_{i_{l}}^{k})\}$, $l=1,2,\ldots,N_{(k,Q_{k})}$. We write
$\mathcal{S}_{(k,Q_{k})}=\left\{y=y(i_{1},\ldots,i_{N_{(k,Q_{k})}}):(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}\right\}$
and $c_{k,Q_{k}}=n_{k}\cdot N_{(k,Q_{k})}+(N_{(k,Q_{k})}-1)m_{(k,Q_{k})}$. Since $\mathcal{C}_{(k,Q_{k})}$ is an $(n_{k},4\epsilon^{\ast})$-separated set, this implies that $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set. Then, by (3.1.2) and (3.9), one has the estimation for the cardinality of $\mathcal{S}_{(k,Q_{k})}$:
@@
Similar to the arguments in (3.8) one has
$D(\mathcal{E}_{c_{k,j}}(y_{k,j}),\mu_{(k,j)})<\frac{2}{k}+\delta_{(k,j)},\ \forall\ y_{k,j}\in\mathcal{S}_{(k,j)}~{}\text{with}~{}k\geq 1\ \text{and }1\leq j<Q_{k},$
@@
Hence, noting that $\mu_{(k,Q_{k})}=\mu^{*}$, in an unified manner we always have
@@
Step $2$. Construct the $\mathbb{F}_{(k,j)}$ and the Moran set $\mathbb{F}$.
One may regard this process as an analogue of the construction of the middle three Cantor set of the unit interval $[0,1]$.
Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$. Suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$, and then we will describe how to construct $\mathcal{I}_{(k,j)}$. Let
@@
#$\#\mathcal{S}_{(k,Q_{k})}\geq\exp\left[(M_{(k,Q_{k})}-M_{(k-1,Q_{k-1})})(1-\zeta)^{2}(H-3\zeta)\right].$ (3.12)
For $x\in\mathcal{I}_{(k,j)^{\ast}}$ and $y\in\mathcal{S}_{(k,j)}$, by the specification property let $z=z(x,y)\in X$ be some point such that
@@
We set $\mathcal{I}_{(k,j)}=\{z(x,y):x\in\mathcal{I}_{(k,j)^{\ast}},y\in\mathcal{S}_{(k,j)}\}$. For $x\in\mathcal{I}_{(k,Q_{k})^{\ast}}$ and two distinct $y,y^{\prime}\in\mathcal{S}_{(k,Q_{k})}$, since $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set, one has
$d_{M_{(k,Q_{k})}}(z(x,y),z(x,y^{\prime}))>2\epsilon^{\ast},$
@@
#$\#\mathcal{I}_{(k,j)}=\#\mathcal{I}_{(k,j)^{\ast}}\cdot\#\mathcal{S}_{(k,j)}=\Pi_{(k^{\prime},j^{\prime})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},j^{\prime})}=\Pi_{(k^{\prime},Q_{k^{\prime}})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}.$ (3.13)
@@
Then for each $(k,j)\in\mathcal{A}$, the following statements are satisfied:
@@
$(2)$ if $z\in\mathcal{I}_{(k,j)}$ descends from $x\in\mathcal{I}_{(k,j)^{\ast}}$, one has
$\overline{B}_{M_{(k,j)}}(z,2\epsilon_{(k,j)})\subset\overline{B}_{M_{(k,j)^{\ast}}}(x,2\epsilon_{(k,j)^{\ast}}).$
This implies that $\{\mathbb{F}_{(k,j)}:(k,j)\in\mathcal{A}\}$ is a sequence of decreasing nonempty closed subsets of $X$. Let
@@
The fractal set $\mathbb{F}$ is contained in $G_{K}^{\omega,U}$, i.e.,
@@
Proof of Lemma 3.4.
By the construction of $\mathbb{F}$, we know
$\mathbb{F}\subset\mathbb{F}_{(0,0)}=\overline{B}_{n(0,0)}(a_{0},\frac{\epsilon^{\ast}}{8})\subset\overline{B(a_{0},\tilde{\epsilon})}\subset U.$
For $p\in\mathbb{F}$ and any $k\geq 1$, we have $p\in\mathbb{F}_{(k,0)}$. Hence, for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there exists $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<3\epsilon_{(k,0)}$. This implies that the orbit of $p$ is $4\epsilon_{(k,0)}$-dense in $X$. By the arbitrariness of $k\geq 1$ and $p\in\mathbb{F}$, we have
$\mathbb{F}\subset\{x\in X:\omega_{f}(x)=X\}.$
For $p\in\mathbb{F}$ and sufficiently large $n\in\mathbb{N}$, to estimate $D(\mathcal{E}_{n}(p),K)$ we split into four cases to discuss.
@@
Case $(2)$. There exists $(k,j)\in\mathcal{A}$ with $j\neq 0$ such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
@@
Case $(3)$. There exist $(k,1)\in\mathcal{A}$ and $i>1$ such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
@@
Case $(4)$. There exist $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$ such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
@@
To sum up, whatever case we consider, by letting $n\to\infty$ we get $V_{f}(p)\subset K$. Based on the above arguments, one can also show that the sequences $\{\mathcal{E}_{n}(p)\}_{n\in\mathbb{N}}$ and $\{\mu_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ have the same limit points. Notice that $K=\overline{\{\mu_{(k^{\prime},j^{\prime})}:(k^{\prime},j^{\prime})>(k,j)\}}$ holds for any fixed $(k,j)\in\mathcal{A}$. Consequently, we have $V_{f}(p)=K$, and hence $\mathbb{F}\subset G_{K}$. ∎
Step $3$. Construct the measure $\rho$ supported on $\mathbb{F}$.
Now we define a probability measure $\rho$ on $\mathbb{F}$ satisfying the $\rho$-measure of some appropriate dynamical balls decays exponentially fast along the sequence $\{M_{k}\}$, where $M_{k}:=M_{(k,Q_{k})}$. Then we can invoke the Entropy Distribution Principle developed in Lemma 2.2 to estimate the lower bound of $\epsilon$-packing topological entropy of $\mathbb{F}$.
@@
For simplicity, we write $\mathcal{I}_{k}$, $\mathbb{F}_{k}$ instead of $\mathcal{I}_{(k,Q_{k})}$, $\mathbb{F}_{(k,Q_{k})}$ respectively. Since $M(X)$ is compact, we may assume that $\rho_{k}\to\rho\in M(X)$ as $k\to\infty$. It is clear that $\rho_{k}(\mathbb{F}_{k})=1$ for every $k$, and for any $k>k_{0}$, one has $\rho_{k}(\mathbb{F}_{k_{0}})=1$ since $\mathbb{F}_{k}\subset\mathbb{F}_{k_{0}}$. Thus one has $\rho(\mathbb{F}_{k_{0}})\geq\limsup_{k\to\infty}\rho_{k}(\mathbb{F}_{k_{0}})=1$. This implies that $\rho(\mathbb{F})=1$. Recall that $M_{(0,0)}=n_{(0,0)}=1$, by (3.12) and (3.13), we have
#$\#\mathcal{I}_{k}\geq\exp\left[(M_{k}-1)\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
@@
Let $k_{0}\geq 1$, $z\in X$ and $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ be any Bowen ball that intersects $\mathbb{F}$. Then for any $k>k_{0}$, one has
@@
Proof of Lemma 3.5.
Since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}\not=\emptyset$, the set $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}_{k_{0}}$ is also nonempty by the construction of $\mathbb{F}$ and #$\#(\mathbb{F}_{k_{0}}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))=1$. Notice that $\mathbb{F}_{k_{0}+1}$ descends from $\mathbb{F}_{k_{0}}$. Then one has
@@
$\rho_{k}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{\Pi_{k_{0}+1\leq k^{\prime}\leq k}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}}{\#\mathcal{I}_{k}}=\frac{1}{\#\mathcal{I}_{k_{0}}}.$
For sufficiently large $k_{0}$ and any Bowen ball $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ with $z\in\mathbb{F}$, since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ is open, one has
@@
$\leq\exp\left[-(M_{k_{0}}-1)(1-\zeta)^{2}\cdot(H-3\zeta)\right]$
$\leq\exp\left[-M_{k_{0}}(1-\zeta)^{3}\cdot(H-3\zeta)\right]$
@@
$(1-\zeta)^{3}\cdot(H-3\zeta)\leq h_{top}^{P}(f,\mathbb{F},\frac{\epsilon^{\ast}}{12})\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$
@@
Dividing in both sides of the above inequality by $|\log\epsilon^{\ast}|$, and then taking $\limsup_{\epsilon^{\ast}\to 0}$, we have
@@
$\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
which proves Theorem 1.1.